# Progressive Token Length Scaling in Transformer Encoders for Efficient Universal Segmentation

### v1

@@ Abstract
A powerful architecture for universal segmentation relies on transformers that encode multi-scale image features and decode object queries into mask predictions. With efficiency being a high priority for scaling such models, we observed that the state-of-the-art method Mask2Former uses $>$50% of its compute only on the transformer encoder. This is due to the retention of a full-length token-level representation of all backbone feature scales at each encoder layer. With this observation, we propose a strategy termed PROgressive Token Length SCALing for Efficient transformer encoders (PRO-SCALE) that can be plugged-in to the Mask2Former-style segmentation architectures to significantly reduce the computational cost. The underlying principle of PRO-SCALE is: progressively scale the length of the tokens with the layers of the encoder. This allows PRO-SCALE to reduce computations by a large margin with minimal sacrifice in performance ($\sim$52GFLOPs reduction with no drop in performance on COCO dataset). We validate our framework on multiple public benchmarks.
@@ 1 Introduction
The tasks of image segmentation (instance [1], semantic [2], and panoptic [3]) are recently being addressed together under the paradigm of “universal” image segmentation [4, 5, 6, 7, 8, 9]. This is due to the evolution of transformer-based [10] approaches that can represent both stuff and things categories [3]) using general tokens, leading to a diminished distinction among the tasks of semantic, instance, and panoptic segmentation. Success of the state-of-the-art universal segmentation framework Mask2Former (M2F) [5] can be attributed to its DEtection TRansformer [11] or DETR-style architecture. This DETR-style segmentation (henceforth termed as M2F-style) framework exhibit exceptional performance across various segmentation tasks without the need for task-specific design choices, setting them apart from preceding modern panoptic segmentation frameworks [12, 13, 14, 15].
@@
The strong performance of M2F-style architecture, however, incurs significant computational overhead hindering their scalability. In this paper, we address this important problem of designing an efficient M2F-style architecture for universal segmentation model. In particular, we present PROgressive Token Length SCALing for Efficient transformer encoders or PRO-SCALE that reduces the computational load occurring in the transformer encoder of such models with surprisingly low performance deterioration.
@@
Figure 1: Compute distribution. (a)) Mask2Former-style segmentation model (b) In the Mask2Former model using Res50 [16] and SWIN-T [17] backbones, the transformer encoder contributes the most to the overall computation cost, accounting for 54.04% and 50.38%, respectively.
@@
M2F-style architectures contain a backbone (for multi-scale feature extraction from input images), a transformer encoder (to capture long-range dependencies and contextual relationships across the multi-scale backbone features), and a transformer decoder (for predicting the masks and labels) along with a segmentation module (see Fig. 1). In efficient segmentation models, where the backbones are traditionally lightweight (SWIN-T [17] and Res50 [16]), cross-scale feature attention has been shown as a potent way to achieve high segmentation performance [5, 18]. However, it has been shown in [18, 19, 20] that due to a large number of query tokens introduced from multi-scale features of light backbones, the encoder incurs the highest computational cost. We make a similar observation in Fig. 1 for M2F where more than 50% of the compute comes from the encoder, given the standard backbones SWIN-T and Res50.
@@
Existing solutions [18, 19] for making DETR-style architectures efficient are only designed to handle detection tasks. For example, Lite-DETR [18] proposed to update larger and smaller scale features in different frequencies for efficient computation. RT-DETR [19] on the other hand, proposed a hybrid encoder that transforms multi-scale features into a sequence of image features through intra-scale interaction and cross-scale fusion. Such detection-based strategies do not suit the segmentation task: either they propose to discard the use of multi-scale features (as in [19]) or lack the ability to reduce computations induced due to constructing a pixel embedding map (as in [18]). This significantly reduces their impact on performance-efficiency trade-off for segmentation models (e.g. $\sim$33GFLOPs reduction but 11% segmentation accuracy drop).
@@
To this end, we introduce PRO-SCALE, an efficient transformer encoder that has three key properties: (1) progressively increase the token length or input size at each encoder layer by introducing larger scale features in the deeper layers, (2) enhance small-scale features by explicitly modeling their inter-dependencies with the channels of higher scale features using a Token Re-Calibration (TRC) module, (3) simplifying the pixel embedding layer by replacing it with a Light Pixel Embedding (LPE) module. The fundamental idea of PRO-SCALE is to address the possible redundancy that arises from consistently maintaining a full-length token sequence across all layers in the encoder. For example as shown in Fig. 2, M2F uses tokens from multi-scale features across all encoder layers resulting in expensive computations. With PRO-SCALE’s progressively expanding tokens, the reduction in sequence length leads to significant FLOPs savings with minimal to negligible degradation in segmentation performance (see Fig. 2). Extensive experiments show that PRO-SCALE based M2F-style architecture achieves a $\sim$52reduction in transformer encoder GFLOPs while maintaining the same segmentation performance. In particular, PRO-SCALE-M2F with SWIN-T backbone achieves a 52.82% PQ with 171.7 GFLOPs (vs. original performance of 52.03% PQ with 234.5 GFLOPs) on the COCO [21] dataset. To summarize, we make the following contributions:
  • We present an efficient transformer encoder PRO-SCALE for M2F-style universal segmentation.
@@
  • PRO-SCALE operates on the fundamental idea of progressively expanding tokens along the encoder depth to address the computation redundancy. It is further assisted by a token re-calibration module and light pixel embedding module that effectively target performance improvement and computational reduction, respectively.
  • Extensive experiments show that PRO-SCALE achieves the best performance-efficiency trade-off on two public benchmarks across diverse settings.
@@ 2 Related Works
Figure 2: Key idea and performance comparison of PRO-SCALE w.r.t. prior works. (a) illustrates the key idea of PRO-SCALE to progressively extend the token length in the transformer encoder. $\{{\bm{s}}_{2},{\bm{s}}_{3},{\bm{s}}_{4}\}$ represent different resolutions. In (b), we show two instantiates of our proposed transformer encoder PRO-SCALE, compared with Mask2Former (M2F) [5] and RT-M2F (an adaptation of [19]). PRO-SCALE eliminates 80.43% and  52% of encoder GFLOPs from M2F while maintaining the competitive performance. Results are computed on the COCO [21] dataset.
DETR [11] embraces an end-to-end object detection approach with a set-prediction objective, discarding the need for manually crafted modules like anchor design and non-maximum suppression. Using the set prediction mechanism introduced in DETR, MaskFormer [4] proposed an approach that converts any existing per-pixel classification model into a mask classification model. This resulted in a universal segmentation architecture that demonstrated state-of-the-art performance across all segmentation tasks on public benchmarks in diverse settings [5, 22, 23, 8, 9] without task-specific design choices. In particular, it employs a transformer decoder to predict a set of pairs (a class prediction, a binary mask). Mask2Former [5] improved MaskFormer by using masked attention in the transformer decoder to restrict the attention to localized features centered around predicted segments. Recently, [22] presented a DETR-style multi-task architecture by extending DINO [24] for both detection and segmentation tasks. PEM [8] introduced a MaskFormer based model that includes a multi-scale feature pyramid network to extract high-semantic-content features with context-based self-modulation to ensure efficiency. PRO-SCALE focuses on making Mask2Former-style universal segmentation models efficient.
In recent times, there have been some interesting efforts [25, 26, 27, 28, 15, 29, 8, 30, 9] to make task-specific or non-M2F-style panoptic segmentation models efficient. For example, YOSO [13] introduced a lightweight panoptic segmentation model by utilizing a feature pyramid aggregator and separable dynamic decoders. ReMax [15] proposed a training pipeline to address the dominant impact of false positive mask assignments in panoptic segmentation by utilizing a separate semantic prediction head. Above methods contain specific model design choices or training strategies suited for the panoptic segmentation task. Unlike these, PRO-SCALE does not conform to one segmentation task and still shows competitive performances.
@@ 3 Methodology
Following Mask2Former (M2F), PRO-SCALE is incorporated in a M2F-style framework. It is composed of a lightweight backbone, our novel transformer encoder design PRO-SCALE, and a transformer decoder with map and class prediction heads (each described next). The overall model framework is shown in Fig. 3. In particular, the input image is fed to the backbone to create multi-scale features. These features are flattened and updated by PRO-SCALE (introduced in Sec. 3.1) with gradual expansion of input tokens along with depth of intermediate encoder layers, allowing a better performance-efficiency trade-off. We further introduce a Token Re-calibration module (described in Sec. 3.1.1) within PRO-SCALE that enriches small-scale features with high-scale features to gain segmentation performance without significant computational overhead. The transformer decoder uses the representations from PRO-SCALE and learnable object queries to compute mask embeddings. Finally, a segmentation module uses the decoder output and per-pixel embeddings from PRO-SCALE for mask predictions. PRO-SCALE provides these per-pixel embeddings required for mask prediction using a parameter-free Light-Pixel Embedding module (described in Sec. 3.1.2). We start by introducing notations and model components, before diving into the details.
@@
  • A backbone network that extracts features from an input image ${\bm{I}}\in\mathbb{R}^{H\times W\times 3}$. The backbone network can provide multi-scale feature maps $\{\bar{{\bm{s}}}_{1},\bar{{\bm{s}}}_{2},\bar{{\bm{s}}}_{3},\bar{{\bm{s}}}_{4}\}$. The spatial resolutions are typically $\nicefrac{{1}}{{4^{2}}}$, $\nicefrac{{1}}{{8^{2}}}$, $\nicefrac{{1}}{{16^{2}}}$, and $\nicefrac{{1}}{{32^{2}}}$ of the input image, respectively. We will denote the token form of these features as ${\bm{s}}_{1},{\bm{s}}_{2},{\bm{s}}_{3}$, and ${\bm{s}}_{4}$, respectively.
@@
  • A Transformer encoder, that enhances the image features $\{{\bm{s}}_{2},{\bm{s}}_{3},{\bm{s}}_{4}\}$ for the transformer decoder and creates a per-pixel embedding map from ${\bm{s}}_{1}$ for the final segmentation module. Feature enhancement is performed on a token representation ${\bm{P}}\in\mathbb{R}^{K\times C}$, where $K=\sum(\nicefrac{{HW}}{{32^{2}}}+\nicefrac{{HW}}{{16^{2}}}+\nicefrac{{HW}}{{8^{2}}})$ and $C=256$ [5]. This ${\bm{P}}$ is composed of concatenated $\{{\bm{s}}_{2},{\bm{s}}_{3},{\bm{s}}_{4}\}$ obtained via flattening the spatial dimensions. The transformer encoder usually consists of several stacked transformer blocks. Following M2F, our framework also consists of six deformable attention transformer [31] layers that includes a self-attention block and a feed-forward block.
@@
  • A Transformer decoder along with segmentation module that decodes binary masks from the modulated image features using a set of randomly initialized object queries. The transformer decoder consists of three stages, with each stage consisting of three transformer layers [10]. Each layer includes a self-attention block, a cross-attention block, and a feed-forward block. Each layer only handles tokens of one scale of features for efficiency [5].
@@
As described above, the original transformer encoder results in significantly expensive computations as it maintains the complete token length of ${\bm{P}}$ for all layers. Simply dropping larger scale features results in poor localization, especially in small objects [18]. For example, removing features ${\bm{s}}_{1}$ and ${\bm{s}}_{2}$ results in a degradation of $\sim 5\%$ AP compared to original model. Moreover, large-scale features are required to produce pixel embedding maps for strong segmentation performance. We will now explain how our approach tackles these obstacles.
@@
Figure 3: Proposed framework. Our model includes our transformer encoder PRO-SCALE (Sec. 3.1), designed to reduce the computational load. $\{{\bm{s}}_{i}\}$s represent the multi-scale backbone features. PRO-SCALE uses a Token Re-Calibration (TRC) module (Sec. 3.1.1) to enhance $s_{4}$ and $s_{3}$ using $s_{2}$ and outputs $\widehat{s}_{4}$ and $\widehat{s}_{3}$, respectively. Further, $s_{1}$ goes through our efficient Light-Pixel Embedding (LPE) module (Sec. 3.1.2) to create pixel embeddings for mask prediction. $p_{1}$, $p_{2}$ and $p_{3}$ represent encoder layer frequency. $\{s^{\prime},s^{\prime\prime},s^{\prime\prime\prime}\}$ represent the outputs of respective layers.

### v2

@@ Abstract
A powerful architecture for universal segmentation relies on transformers that encode multi-scale image features and decode object queries into mask predictions. With efficiency being a high priority for scaling such models, we observed that the state-of-the-art method Mask2Former uses $>$50% of its compute only on the transformer encoder. This is due to the retention of a full-length token-level representation of all backbone feature scales at each encoder layer. With this observation, we propose a strategy termed PROgressive Token Length SCALing for Efficient transformer encoders (PRO-SCALE) that can be plugged-in to the Mask2Former segmentation architecture to significantly reduce the computational cost. The underlying principle of PRO-SCALE is: progressively scale the length of the tokens with the layers of the encoder. This allows PRO-SCALE to reduce computations by a large margin with minimal sacrifice in performance ($\sim$52encoder and $\sim$ 27% overall GFLOPs reduction with no drop in performance on COCO dataset). Experiments conducted on public benchmarks demonstrates PRO-SCALE’s flexibility in architectural configurations, and exhibits potential for extension beyond the settings of segmentation tasks to encompass object detection. https://github.com/abhishekaich27/proscale-pytorch
@@ 1 Introduction
The tasks of image segmentation (instance (He et al., 2017), semantic (Tu, 2008), and panoptic (Kirillov et al., 2019)) are recently being addressed together under the paradigm of “universal” image segmentation (Cheng et al., 2021; 2022; Jain et al., 2023; Gu et al., 2024; Cavagnero et al., 2024; Rosi et al., 2024). This is due to the evolution of transformer-based (Vaswani et al., 2017) approaches that can represent both stuff and things categories (Kirillov et al., 2019)) using general tokens, leading to a diminished distinction among the tasks of semantic, instance, and panoptic segmentation. Success of the state-of-the-art universal segmentation framework Mask2Former (M2F) (Cheng et al., 2022) can be attributed to its DEtection TRansformer (Carion et al., 2020) or DETR-style architecture. This DETR-style segmentation (henceforth termed as M2F-style) framework exhibits exceptional performance across various segmentation tasks without the need for task-specific design choices, setting them apart from preceding modern panoptic segmentation frameworks (Rashwan et al., 2024; Hu et al., 2023; Ammar et al., 2023; Sun et al., 2023).
@@
The strong performance of M2F-style architecture, however, incurs significant computational overhead hindering their widespread deployment. In this paper, we address this important problem of designing an efficient M2F-style architecture for universal segmentation model. In particular, we present PROgressive Token Length SCALing for Efficient transformer encoders or PRO-SCALE that reduces the computational load occurring in the transformer encoder of such models with surprisingly low performance deterioration.
@@
Figure 1: Compute distribution. (a) Mask2Former-style segmentation model (b) In the Mask2Former model using Res50 (He et al., 2016) and SWIN-T (Liu et al., 2021) backbones, the transformer encoder contributes the most to the overall computation cost, accounting for 54.04% and 50.38%, respectively.
@@
M2F-style architectures contain a backbone (for multi-scale feature extraction from input images), a pixel decoder or transformer encoder (to capture long-range dependencies and contextual relationships across the multi-scale backbone features), and a transformer decoder (for predicting the masks and labels) along with a segmentation module (see Fig. 1). In efficient segmentation models, where the backbones are traditionally lightweight (SWIN-T (Liu et al., 2021) and Res50 (He et al., 2016)), cross-scale feature attention has been shown as a potent way to achieve high segmentation performance (Cheng et al., 2022; Li et al., 2023a). However, it has been shown in (Li et al., 2023a; Lv et al., 2023; Yao et al., 2021) that due to a large number of query tokens introduced from multi-scale features of light backbones, the encoder incurs the highest computational cost. We make a similar observation in Fig. 1 for M2F where more than 50% of the compute comes from the encoder, given backbones SWIN-T and Res50.
@@
Existing solutions (Li et al., 2023a; Lv et al., 2023) for making DETR-style architectures efficient are only designed to handle detection tasks. For example, Lite-DETR (Li et al., 2023a) proposed to update larger and smaller scale features in different frequencies for efficient computation. RT-DETR (Lv et al., 2023) on the other hand, proposed a hybrid encoder that transforms multi-scale features into a sequence of image features through intra-scale interaction and cross-scale fusion. Such detection-based strategies do not suit the segmentation task: either they propose to discard the use of multi-scale features (as in (Lv et al., 2023)) or lack the ability to reduce computations induced due to constructing a pixel embedding map (as in (Li et al., 2023a)). This significantly reduces their impact on performance-efficiency trade-off for segmentation models (e.g. $\sim$33GFLOPs reduction but 11% segmentation accuracy drop).
@@
To this end, we introduce PRO-SCALE for M2F’s transformer encoder that has two key properties: (1) progressively increase the token length or input size at each encoder layer by introducing larger scale features in the deeper layers (2) simplifying the pixel embedding layer by replacing it with a Light Pixel Embedding (LPE) module. The fundamental idea of PRO-SCALE is to address the possible redundancy that arises from consistently maintaining a full-length token sequence across all layers in the encoder. For example as shown in Fig. 2, M2F uses tokens from multi-scale features across all encoder layers resulting in expensive computations. With PRO-SCALE’s progressively expanding tokens, the reduction in sequence length leads to significant FLOPs savings with minimal to negligible degradation in segmentation performance (see Fig. 2). Extensive experiments show that PRO-SCALE based M2F architecture achieves a $\sim$52reduction in transformer encoder GFLOPs while maintaining the same segmentation performance. In particular, PRO-SCALE-M2F with SWIN-T backbone achieves a 52.82% PQ with 171.7 GFLOPs (vs. original performance of 52.03% PQ with 234.5 GFLOPs) on the COCO (Lin et al., 2014) dataset.
@@
To summarize, we present an efficient transformer encoder PRO-SCALE for M2F universal segmentation architecture. PRO-SCALE operates on the fundamental idea of progressively expanding tokens along the encoder depth to address the computation redundancy. It is further assisted by a light pixel embedding module that effectively target computational reduction in the encoder module. Extensive experiments show that PRO-SCALE achieves the best performance-efficiency trade-off on two datasets across diverse settings.
@@ 2 Related Works
Figure 2: Key idea and performance comparison of PRO-SCALE w.r.t. prior works. (a) illustrates the key idea of PRO-SCALE to progressively extend the token length in the transformer encoder. $\{{\mathbf{s}}_{2},{\mathbf{s}}_{3},{\mathbf{s}}_{4}\}$ represent different resolutions. In (b), we show two instantiates of our proposed transformer encoder PRO-SCALE, compared with Mask2Former (M2F) (Cheng et al., 2022) and RT-M2F (an adaptation of (Lv et al., 2023)). PRO-SCALE eliminates 80.43% (with configuration ($p_{1}$, $p_{2}$, $p_{3}$) = (1,1,1)) and 51.98% (with configuration ($p_{1}$, $p_{2}$, $p_{3}$) = (3,3,3)) of encoder GFLOPs from M2F while maintaining the competitive performance. Results are computed on the COCO (Lin et al., 2014) dataset.
DETR (Carion et al., 2020) embraces an end-to-end object detection approach with a set-prediction objective, discarding the need for manually crafted modules like anchor design and non-maximum suppression. Using the set prediction mechanism introduced in DETR, MaskFormer (Cheng et al., 2021) proposed an approach that converts any existing per-pixel classification model into a mask classification model. This resulted in a universal segmentation architecture that demonstrated state-of-the-art performance across all segmentation tasks on public benchmarks in diverse settings (Cheng et al., 2022; Li et al., 2023b; Ding et al., 2022; Cavagnero et al., 2024; Rosi et al., 2024) without task-specific design choices. In particular, it employs a transformer decoder to predict a set of pairs (a class prediction, a binary mask). Mask2Former (Cheng et al., 2022) improved MaskFormer by using masked attention in the transformer decoder to restrict the attention to localized features centered around predicted segments. Recently, (Li et al., 2023b) presented a DETR-style multi-task architecture by extending DINO (Zhang et al., 2022) for both detection and segmentation tasks. PEM (Cavagnero et al., 2024) introduced a MaskFormer based model that includes a multi-scale feature pyramid network to extract high-semantic-content features with context-based self-modulation to ensure efficiency. PRO-SCALE focuses on making the Mask2Former universal segmentation model efficient.
In recent times, there have been some interesting efforts (Wang et al., 2019; de Geus et al., 2020; Hong et al., 2021; Hou et al., 2020; Sun et al., 2023; Šarić et al., 2023; Cavagnero et al., 2024; Xu et al., 2024; Rosi et al., 2024) to make task-specific or non-M2F-style panoptic segmentation models efficient. For example, YOSO (Hu et al., 2023) introduced a lightweight panoptic segmentation model by utilizing a feature pyramid aggregator and separable dynamic decoders. ReMaX (Sun et al., 2023) proposed a training pipeline to address the dominant impact of false positive mask assignments in panoptic segmentation by utilizing a separate semantic prediction head. Above methods contain specific model design choices or training strategies suited for the panoptic segmentation task. Unlike these, PRO-SCALE does not conform to one segmentation task and still shows competitive performances.
@@ 3 Methodology
Following M2F, our proposed framework is composed of a lightweight backbone, our novel transformer encoder design PRO-SCALE, and a transformer decoder with mask and class prediction heads (each described next). The overall model framework is shown in Fig. 3. In particular, the input image is fed to the backbone to create multi-scale features. These features are flattened and updated by PRO-SCALE (introduced in Sec. 3.1) with gradual expansion of input tokens along with depth of intermediate encoder layers, allowing a better performance-efficiency trade-off. The transformer decoder uses the representations from PRO-SCALE and learnable object queries to compute mask embeddings. Finally, a segmentation module uses the decoder output and per-pixel embeddings from PRO-SCALE for mask predictions. PRO-SCALE provides these per-pixel embeddings required for mask prediction using a parameter-free Light-Pixel Embedding module (described in Sec. 3.1.1). We start by introducing notations and model components, before diving into the details.
@@
  • A backbone network that extracts features from an input image ${\mathbf{I}}\in\mathbb{R}^{H\times W\times 3}$. The backbone network can provide multi-scale feature maps $\{\bar{{\mathbf{s}}}_{1},\bar{{\mathbf{s}}}_{2},\bar{{\mathbf{s}}}_{3},\bar{{\mathbf{s}}}_{4}\}$. The spatial resolutions are typically $\nicefrac{{1}}{{4^{2}}}$, $\nicefrac{{1}}{{8^{2}}}$, $\nicefrac{{1}}{{16^{2}}}$, and $\nicefrac{{1}}{{32^{2}}}$ of the input image, respectively. We will denote the token form of these features as ${\mathbf{s}}_{1},{\mathbf{s}}_{2},{\mathbf{s}}_{3}$, and ${\mathbf{s}}_{4}$, respectively.
@@
  • A pixel decoder or transformer encoder, that enhances the image features $\{{\mathbf{s}}_{2},{\mathbf{s}}_{3},{\mathbf{s}}_{4}\}$ and also creates a per-pixel embedding map from ${\mathbf{s}}_{1}$ and ${\mathbf{s}}_{2}$. The feature enhancement is performed on a token representation ${\mathbf{P}}\in\mathbb{R}^{K\times C}$, where $K=\sum(\nicefrac{{HW}}{{32^{2}}}+\nicefrac{{HW}}{{16^{2}}}+\nicefrac{{HW}}{{8^{2}}})$ and $C=256$ (Cheng et al., 2022). This ${\mathbf{P}}$ is composed of concatenated $\{{\mathbf{s}}_{2},{\mathbf{s}}_{3},{\mathbf{s}}_{4}\}$ obtained via flattening the spatial dimensions and fed to a transformer encoder. The transformer encoder usually consists of several stacked transformer blocks. Following M2F, our framework also consists of six deformable attention transformer (Zhu et al., 2020) layers that includes a self-attention block and a feed-forward block. The per-pixel embedding map, on the other hand, is generated using convolutional layers ${\mathbf{s}}_{1}$ and ${\mathbf{s}}_{2}$, which enhance local spatial details to be used to create the segmentation mask.
@@
  • A Transformer decoder along with segmentation module that decodes binary masks from the modulated image features from the pixel decoder using a set of randomly initialized object queries. The transformer decoder consists of three stages, with each stage consisting of three transformer layers (Vaswani et al., 2017). Each layer includes a self-attention block, a cross-attention block, and a feed-forward block. Each layer only handles tokens of one scale of features for efficiency (Cheng et al., 2022).
@@
As described above, the original transformer encoder results in significantly expensive computations from global attention mechanism, as it maintains the complete token length of ${\mathbf{P}}$ for all layers. Simply dropping larger scale features results in poor localization, especially in small objects (Li et al., 2023a). For example, removing features ${\mathbf{s}}_{1}$ and ${\mathbf{s}}_{2}$ results in a degradation of $\sim 5\%$ AP compared to original model. Moreover, the use of large-scale features to produce pixel embedding map in the convolutional layers require processing an enormous number of tokens, leading to an extremely high computational demand. We will now explain how our approach tackles these obstacles.
@@
Figure 3: Proposed framework. Our model includes our transformer encoder PRO-SCALE (Sec. 3.1), designed to reduce the computational load. $\{{\mathbf{s}}_{i}\}$s represent the multi-scale backbone features. PRO-SCALE progressively scale the length of the tokens with the layers of the encoder. This allows PRO-SCALE to reduce computations by a large margin with minimal sacrifice in performance. Further, $s_{1}$ goes through our efficient Light-Pixel Embedding (LPE) module (Sec. 3.1.1) to create pixel embeddings for mask prediction. $p_{1}$, $p_{2}$ and $p_{3}$ represent encoder layer frequency. $\{s^{\prime},s^{\prime\prime},s^{\prime\prime\prime}\}$ represent the outputs of respective layers.